OpenStudy (anonymous):

Solve the following system using the substitution method. 3x – y = –2 6x + 5y = 24

OpenStudy (anonymous):

x=2/3 you got it right?

OpenStudy (anonymous):

from equation 1 y = 3x+2 plug this in to 2: 6x + 5 (3x+2) = 24 21x + 10 = 24 x = 14/21 = 2/3 pkug 2/3 into ist equn: 3(2/3) - y = -2 -y = -2-2 = -4 y = 4 solution: x = 2/3 , y = 4